Solve the following problems by providing the mathematical and pictorial representations:

1. Base problem (by G Lavery)

Baseballs are manufactured in groups of 4, sporting goods stores sell the baseballs in packages of 3. Therefore, 3/4 of a group from the factory are used to make a store package. How many store packages can be made from 3 factory groups of baseballs?

2. Cookies trays problem (by Pollino)

Exactly 7 cookies fit on each tray for baking four - sevenths of a tray is used to make a small box of cookies. If a baker has made8 trays of cookies, how many boxes can be filled?


1
Expert's answer
2021-06-22T17:34:55-0400

1:



Then, 1 group from the factory is used to make 4÷3 store packages.

{Simply by using the unitary method,}

Therefore, 3 factory groups are used to make "\frac{4\times3}{3}"  =4 store packages.


2:

Given,



No. Of cookies fit in each tray= 7

four - sevenths of a tray is used to make a small box of cookies.

Total number of trays= 8

According to given,

Number of cookies in a box= 7×(4÷7)= 4 cookies

Total number of cookies made= Number of trays* Number of cookies in each tray= 8×7= 56

Therefore,

Total number of boxes= Total number of cookies/Number of cookies in each box

= 56÷4= 14

Hence,

A total of 14 boxes can be filled.